OpenStudy (anonymous):

Explain, in complete sentences, how you would use the graphing method to solve the following system of inequalities and what the final solution will look like on the coordinate plane. x – 2y –8 7x + y < –6

OpenStudy (anonymous):

x – 2y <–8 [I added the inequality as it did not make sense] 7x + y < –6 convert to equations temporarily x – 2y =–8 7x + y = –6 In both equations set x at 0 to find y-intercept then set y at 0 to find x-intercept x – 2y =–8 (0, 4) and (-8,0) join a dashed line through these points and shade above the line indicating points that abey the inequality. 7x + y = –6 (0,-6) and (-6/7, 0) join a dashed line through these points and shade (in an opposite direction to the first shading) below the line. Where you find cross hatching from your shading indicates the area that satisfies both inequalities. Voila!